Grade 12Mechanics

The coordinate of a particle moving along x-axis varies with time as x= 6- 4t-t² where x in meter and t in second then distance covered by the particle in first 3 seconds is ? (Answer = 5m)

Ayushi
9 Years agoGrade 12

1 Answer

Seemant Sinha
9 Years ago
X= 6-4t +t²Put t=0X=6Put t=1X=3Put t= 2X=2Put t=3X=3So distance traveled In 1st second = 6-3 = 3distance traveled In 2nd second = 3-2 = 1And distance traveled In 3rd second = |2-3| = 1Now add those distances 3+1+1=5m